OpenStudy (anonymous):

Determine if the first polynomial is a factor of the second polynomial: X+1; X^3+X^2-4x-4

OpenStudy (anonymous):

If (x+a) if a factor of f(x) , f(-a) = 0

OpenStudy (amistre64):

you cant argue with that ;)

OpenStudy (anonymous):

more explanation please?

OpenStudy (amistre64):

when two numbers are multiplied together to form a new number; you know that the factors of that new number will never have a remainder after division.

OpenStudy (anonymous):

let f(x) = (x-a)Q(x) + R Where Q(x) is the quotient, and R is the remainder (when divided by (x+a) ) let x = a f(a) = (a - a)(Q(x)) + R = R. If f(a) = 0, then it is a factor. QED, featheres.

OpenStudy (anonymous):

When divided by (x-a)*

OpenStudy (amistre64):

for example, is 8 an integer factor if 15? 15/8 = 1 plus a remainder of 7. so no. is 3 an inteeger factor of 15? 15/3 = 5 with a remainder of 0. 3 IS a factor.

OpenStudy (amistre64):

you can do it the long way by dividing (x+3) into the other; and see if the remainder is 0.

OpenStudy (anonymous):

that some nice explanation ;) tnkx

OpenStudy (amistre64):

;) Newton helped :)